Documentation for AHCAL TestBeam at CERN SPS in May 2017. Combined TestBeam with HGCAL is described here:AHCALandHGCALTestBeamCERN2017

Contents

Setup

May 2017

In May, we have take data in the H2@SPS line at CERN, during the period between the 24/05/2017 to the 31/05/2017.

Our setup consisted of a small steel absorber structure with 17 plates (thickness 19 mm), equipped with 15 active layers of 1 HBU each. The setup was placed within the M1 magnet, and operated with 1.5T B field, and without B field. All layers have individually wrapped tiles. The detector was read out with 2 mini-LDAs, placed outside of the magnet.

The layers were ordered as follows (first is the closest to beam):

Phys.
order

Module#

power
cable

HDMI
cable

MPOD

LDA#

LDA
port#

name

1.

1

01

01

u0,u100,u400

0

00

SM_100

2.

2

02

02

u1,u101,u401

0

01

SM_1

3.

3

03

03

u2,u102,u402

0

02

SM_2

4.

4

04

04

u3,u103,u403

0

03

SM_4

5.

5

05

05

u4,u104,u404

0

05

SM_5

6.

6

06

06

u5,u105,u405

0

06

SM_7

7.

7

16

07

u6,u106,u406

0

07

Mainz_HBU3

8.

8

08

08

u7,u107,u407

0

08

HBU3_1

9.

9

09

09

u200,u300,u500

1

00

HBU3_2

10.

10

10

10

u201,u301,u501

1

01

HBU3_3

11.

11

11

11

u202,u302,u502

1

02

HBU3_4

12.

12

12

12

u203,u303,u503

1

03

HBU3_5

13.

13

13

13

u204,u304,u504

1

06

HBU3_6

14.

14

14

14

u205,u305,u505

1

07

HBU3_7

15.

15

15

15

u206,u306,u506

1

08

HBU3_12

Chip and Channel MAP

attachment:HBU_map_with_beam_info_top_readout.png

MapChipChan2IJK_May2017

mapping.txt

DAQ Setup

The used DAQ hierarchy is as follows:

ASIC (SPIROC 2b, Omega group)

all layers were operating with SPIROC2b

Interfaces

All layers were operating with the new generation of interface boards (CIB, POWER, DIF, CALIB) without magnetic components. For layers 1,8 and 9 (SM100, HBU3_2, HBU3_1) pre-series boards were used (accidentally), where as many magnetic components as possible were removed (see logbook entries of 23 May 2017)

Link Data Agregator (LDA)

Two mini-LDAs were used.

Clock and Control Card (CCC)

The CCC provides the master clock of the DAQ and starts and stops the acquisition according to the spill level and readiness of all ASICs.

It receives the trigger validation signal and the spill signal and distribute them.

The CCC together with the 2 mini-LDAs was placed outside of the magnet in an iron box to shield it from the magnetic field.

Beam Interface (BIF)

Based on AIDA mini-TLU hardware. Was used to get precise time stamps for the trigger signal, and statistics of triggers for the runs

Labview

The Main.vi in f:\CERN_2017_May\LabviewTestbeam2017May\ was used (has its own branch in svn)

EUDAQ

For beam data, EUDAQ2 was used: in /home/calice/Desktop/EUDAQ2-testbeam2017May/ use ./STARTRUN.local (and ./KILLRUN.local)

For LED data, our special branch for EUDAQ1.6 supporting parameter scans was used: in /home/calice/Desktop/EUDAQ1.6/producers/calice/scripts use ./LED_2017 (and ./KILLRUN)

Monitoring

Online monitoring and event display is described in a separate wiki page: http://flcwiki.desy.de/AHCALTestBeamCERN2017/Monitoring

Trigger and spill signal

Trigger

For the validation a coincidence of two CMS owned trigger scintillators of 14*14 cm^2 in ??? m distance in front of the detector was used. This is recorded by the BIF. The coincidence was produced by CMS NIM electronics and provided in the beam hut. From there we put it to the beam area via the patch panel.

Spill

The spill signal was created from the WWE and the EE sipnals provided by SPS. We produced the spill signal in the beam hut (withour NIM electronics) and put it to the beam area via the patch panel.

The spills had ~5s length, and we had 1 or 2 spills per supercycle, where the supercycles were ~38s to ~48s long (depending on LHC filling and AWAKE commissioning). The spill signal was a bit longer than the time we actually got particles, so we have a few readout cycles without triggers at the beginning and end of each spill.

Others

A threshold Cherenkov detector was present in the beamline, but the signal not used. No t0 signals any more, everything now handled by the BIF!

Materials in front of AHCAL

beam area

attachment:materials_in_SPS_H2_beam_line.png

upstream of concrete wall H2_beam_line.pdf

Run Information

run_summary.txt

Database

Pedestal

Folder

DB Tag

Comments

/cd_calice_Ahc2/TestbeamMay2017/Pedestal

ahc2_pedestal_003

Pedestal Muon run, Magnet Off, 20170524 top readout

/cd_calice_Ahc2/TestbeamMay2017/Pedestal

ahc2_pedestal_005

Pedestal Muon run, Magnet On, 20170525 top readout

AHCALTestBeamCERN2017 (last edited 2017-10-18 09:52:28 by yujisudo)